Evaluate:
log ( 7x - 4 ) = 3

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

log ( 7x - 4 ) = 3
7x - 4 = 10^3 = 1,000
7x = 1,004
x = 1,004/7

Answer by stanbon(48569)

Evaluate:
log ( 7x - 4 ) = 3
Rewrite in exponential form:
7x-4 = 10^3
7x-4 = 1000
7x = 1004
x = 143.4286
=======================
Cheers,
Stan H.